# Language-centered Human Activity Recognition

### v2

@@ 3.5.1. Sensor encoder structure
This process is repeated across multiple encoder layers, with the output of each layer serving as the input for the next.
@@ 4.1.1. Dataset
We evaluate the performance of LanHAR using four public datasets that have been extensively employed in existing research (Xu et al., 2023). (1) HHAR (Stisen et al., 2015) includes individuals performing six activities (sitting, standing, walking, going downstairs, going upstairs, and biking) using six different mobile phone types. (2) UCI (Reyes-Ortiz et al., 2016) captures six activities (i.e., standing, sitting, lying, walking, going downstairs, and going upstairs). (3) Motion (Malekzadeh et al., 2019) involves 24 participants performing six activities (sitting, standing, walking, going upstairs, going downstairs, and jogging). (4) Shoaib (Shoaib et al., 2014) includes 10 participants engaged in seven activities (sitting, standing, walking, going upstairs, going downstairs, jogging, and biking).
@@ 4.1.2. Data preprocessing
We preprocess all datasets to ensure consistent sampling rates and window sizes. The sampling rate is set to 20 Hz to simulate a resource-constrained scenario, following prior studies (Xu et al., 2023). For cross-dataset activity recognition, we focus on four common activities (walking, going upstairs, going downstairs, and sitting) across the four datasets. We select one dataset as the source and one of the remaining three as the target, resulting in a total of 12 combinations. Datasets are categorized based on their role as either source or target. For source datasets, each is split into a training set (80%) and a validation set (20%). For target datasets, the entire dataset is used as the test set to directly assess cross-dataset performance. For new activity recognition, we train the model on the four common activities from the source dataset. We then select activities from the target dataset that differ from the four common activities as new activities. Finally, we evaluate the model’s ability to recognize these new activities by determining which activity labels the new data corresponds to.
@@
LanHAR Accuracy 0.835 0.812 0.728 0.860 0.807 0.771 0.693 0.762 0.732 0.723 0.785 0.768 0.773
F1 scores 0.816 0.807 0.694 0.819 0.755 0.712 0.612 0.711 0.720 0.690 0.731 0.723 0.732
4.1.3. Baselines
We compare our method LanHAR with three categories of baselines including conventional HAR, cross-dataset HAR, and new activity baselines.
(i) Conventional HAR baselines: DCNN (Yang et al., 2015) are widely used in the field of human activity recognition for their ability to automatically extract local features from time-series sensor data. Transformers (Vaswani, 2017) utilize self-attention mechanisms to capture long-range dependencies in sensor data.
(ii) Cross-dataset HAR baselines: SDMix (Lu et al., 2022) It is a method for cross-domain activity recognition that uses the mixup method that addresses semantic inconsistencies caused by domain gaps. UDAHAR (Chang et al., 2020) is a method for cross-dataset activity recognition using domain adaptation. UniHAR (Xu et al., 2023) is a method for cross-dataset activity recognition. It designs physical-informed augmentation methods to improve the cross-dataset HAR performance.
(iii) New activity baselines: DCNN (Yang et al., 2015) To adapt it for recognizing new activities, we fine-tune it using a small portion of data from the target dataset. Nuactiv (Cheng et al., 2013) is a method designed for new activities. It predefined activity label-related attributes: actions of a specific part of the body.
(iv) Variants of our model: (1) LanHAR without Subtasks of Text Encoder (w/o STE). We remove the subtasks from the text encoder, focusing its training on alignment. (2) LanHAR without Training of Text Encoder (w/o TTE). We remove the training part of the text encoder and directly use a pre-trained BERT model to replace the text encoder. (3) LanHAR with LLAMA2 (w LLAMA2). We replace GPT-4 with LLAMA2 for obtaining semantic interpretations. (4) LanHAR with GPT-3 (w GPT-3). We replace GPT-4 with GPT-3 for obtaining semantic interpretations. (5) LanHAR with Different Language Model (w DLM). We use the transformer architecture to replace the BERT in LanHAR. (6) LanHAR without Iterative re-Generation (w/o IG). We remove iterative re-generation method. (7) LanHAR without Iterative re-Generation and designed Prompt (w/o IGP). We remove the iterative re-generation method and designed prompt. We directly use a simple prompt only containing data and task requirement.
Table 3. New Activity Performance. Bolded scores indicate the best.
Method Source HHAR UCI Motion Shoaib
DCNN Accuracy 0.197 0.256 0.227 0.243 0.201 0.226 0.212 0.267 0.271 0.189 0.201 0.221
Nuactiv Accuracy 0.304 0.362 0.351 0.467 0.423 0.347 0.425 0.519 0.461 0.372 0.356 0.361
LanHAR Accuracy 0.572 0.556 0.601 0.663 0.721 0.704 0.563 0.586 0.622 0.657 0.661 0.627
4.1.4. Metric
We evaluate our model using two metrics: accuracy and F1 score. For cross-dataset HAR performance, we use accuracy to measure the correctness of the classification results for the selected four activities. For new activities, we use accuracy to evaluate the correctness of the classification results only for the new activity.
4.1.5. Implementation
We implement our method and baselines using PyTorch 1.9.0 in a Python 3.7 environment and train them using two NVIDIA RTX A5000 GPUs, each with 24 GB of memory. For LLMs, we choose GPT4 as our LLMs to obtain semantic interpretations of sensor readings and activity labels. For the text encoder, we use the BERT-base architecture as its structure and initialize it with the pre-trained BERT-base parameters (Devlin et al., 2018). For the sensor encoder, we set the embedding dimension to $768$ for consistency with pre-trained BERT. The multi-head attention parameter is set to $2$ and the encoder layer consists of $3$ layers. We apply the AdamW optimizer with a learning rate of 1e-5 and set the batch size to 32.
4.2. Overall performance
We compare our method with various baselines, with the results presented in Table 2. Overall, our method demonstrates superior average performance across 12 different source and target dataset combinations, leading in the majority of settings. Compared to conventional HAR baselines, our method consistently outperforms them. These traditional baselines are not specifically designed for cross-dataset HAR, and their strong performance typically depends on large, carefully collected datasets. Compared to cross-dataset HAR baselines, our method outperforms the cross-dataset HAR baselines (except in the setting from Shoaib to Motion). Although SDMix and UDAHAR are also designed for cross-dataset HAR, they have certain requirements. For instance, UDAHAR may be more suitable for datasets with smaller distribution gaps. This consistent performance indicates the robustness and reliability of our method in diverse conditions and datasets.
@@ 4.3. New activity performance
We compare our method with two new activity recognition baselines, with the results presented in Table 3. Specifically, we train the model using four common activities from the source dataset and evaluate it on different activities from the target dataset that are not among the four common activities, treating them as new activities. The trained model is then used for recognizing these new activities, i.e., determining which activity new data belong to.
For new activity HAR, our model outperforms the baseline. Our model’s performance is superior to that of a fine-tuned DCNN model, which indicates its advantage over simple fine-tuniing on small datasets. Additionally, our model performs better than Nuactiv, which relies on predefined attributes for activities. Defining these attributes is challenging and varies significantly across different activity contexts, directly impacting Nuactiv’s performance. This limitation hinders Nuactiv’s performance in cross-dataset settings.
@@ 4.4. Category-level HAR performance
We further evaluate the category-level accuracy of HAR performance. An activity is considered correctly classified if it is categorized into a similar activity category, and incorrect if placed into a dissimilar category. Specifically, we group seven activities (walking, jogging, biking, sitting, standing, going upstairs, going downstairs, and lying) into three categories based on their similarity: Category 1 (walking, jogging, biking), Category 2 (sitting, standing, lying), and Category 3 (going upstairs, going downstairs). We compared our method with all the baselines. The average results are shown in Figure 10. As shown, our method improves category-level HAR accuracy compared to the baselines. This improvement has the potential to reduce the impact of classification errors on downstream applications.
@@ 4.5.1. The effect text encoder
To assess the impact of the subtasks in the text encoder, we compare our method with a variant (i.e., w/o STE), as shown in Figure 10. In the w/o STE variant, we remove the subtasks (i.e., the two contrastive learning and reconstruction tasks) and focus only on aligning the semantic interpretations of sensor readings and activity labels. As shown in the figure, removing the subtasks leads to a decline in performance compared to our full model. This demonstrates that the subtasks we designed help the model better understand IMU sensor readings and HAR-related semantics.
To further explore the effect of the text encoder for semantic interpretation alignment, we replace the carefully designed text encoder with a pre-trained BERT model and directly use it for training in stage 2. The results are shown in Figure 10. From the figure, it is clear that the performance of w/o TTE is lower than both w/o STE and our full model. This suggests that pre-trained language models without fine-tuning cannot fully capture HAR-related semantics or effectively align different HAR-related semantic interpretations.
@@
4.5.2. The impact of different pre-trained LM
In our text encoder, we used a pre-trained BERT architecture and fine-tuned it with our designed tasks. To explore the impact of using different language model architectures on HAR performance, we replaced the BERT-based structure with GPT-2. As shown in Figure 12, this replacement resulted in decreased HAR performance. BERT is generally more suitable for tasks that require understanding sentence meaning and comparing text similarities, as its architecture and pre-training are optimized for capturing semantics and contextual relationships in a bidirectional manner. In contrast, GPT-2, with its unidirectional processing, is less effective at understanding sentence-level meaning and comparing text similarity, though it excels in generating coherent and contextually relevant text. This explains the performance drop when GPT-2 is used for HAR tasks.
Figure 11. The impact of different pre-trained LM
Figure 12. The impact of different LLMs
@@ 4.5.3. The impact of different LLMs
Given the varying capabilities of different LLMs, we explore the impact of three LLMs (GPT-4, GPT-3, and LLAMA2) on interpreting sensor readings and activity labels for HAR performance. As shown in Figure 12, the performance of GPT-3 and LLAMA2 is noticeably lower than that of GPT-4, highlighting the importance of the LLM’s capability in language-centered human activity recognition.
4.5.4. The impact of different quality of LLMs’ response
To ensure the quality of the LLM’s interpretation of sensor readings, we (i) design specific prompts to guide the LLM’s responses and (ii) develop an iterative re-generation method to filter out low-quality semantic interpretations and generate improved ones. To demonstrate the effectiveness of the automatic iterative re-generation method, we compared the KL divergence of the same activities across different datasets before and after applying the method. Specifically, for each activity, we measure the KL divergence between any two datasets both with and without applying the iterative re-generation method (i.e., “w iterative re-generation” and “w/o iterative re-generation”), as shown in Figure 14. After applying the method, the average KL divergence for the same activity across different datasets decreased by 58.75%, indicating that the automatic iterative re-generation method significantly improves the quality of LLM responses and reduces the distribution gap in heterogeneous data.
Figure 13. KL divergence with and without the iterative re-generation method
@@ 4.5.4. The impact of different quality of LLMs’ response
To verify the impact of these methods on HAR performance, we compare our method with two variants: one using only simple prompts (w/o IGP) and another using only specifically designed prompts (w/o IG). As shown in Figure 14, the performance when using only simple prompts is the lowest, indicating that LLMs require guidance to generate high-quality responses. The combination of specially designed prompts and the automatic iterative re-generation method performs the best, demonstrating that our approach effectively enhances the quality of LLM responses, leading to improved human activity recognition outcomes.
@@ 4.6. Parameter sensitivity
In our model, selecting the appropriate parameters for the sensor encoder is critical, as it directly affects its suitability for deployment on mobile devices (e.g., ensuring the parameter size is sufficiently small). In this section, we primarily analyze the impact of the embedding dimensions and the number of encoding layers on performance.
@@ 4.6.1. Effect of different number of attention heads
Figure 16 shows the performance comparison $w.r.t.$ the number of attention heads in the sensor encoder. Overall, the performance is relatively stable under different numbers of attention heads. We chose the number of attention heads of 2 because it provides good performance while reducing the model’s parameter size.
@@ 4.6.2. Effect of different number of encoding layers
Figure 16 shows the performance comparison with respect to different numbers of encoding layers in the sensor encoder. As observed, a smaller number of encoding layers (e.g., 2 layers) results in insufficient representation, while a larger number (e.g., 6 layers) increases model complexity. We selected 3 encoding layers as it strikes a balance, delivering good performance without significantly increasing model complexity.
@@
4.7. Complexity analysis
To ensure the sensor encoder is suitable for deployment on mobile devices, we analyze its parameter size and inference time. When deployed on a mobile device, the sensor encoder, along with the activity label embeddings (approximately 0.01 MB of memory), requires around 61 MB of total memory. For inference, the execution time for processing one IMU sample (i.e., recognizing an activity) on a Pixel 7 device is 23.71ms. These results confirm that the sensor encoder can operate efficiently on mobile devices.

### v3

@@ 3.5.1. Sensor encoder structure
This process is repeated across multiple encoder layers, each feeding its output to the next.
@@ 4.1.1. Dataset
We first evaluate the performance of LanHAR using four public datasets with simple activities that have been extensively employed in existing research (Xu et al., 2023). (1) HHAR (Stisen et al., 2015) includes individuals performing six activities (sitting, standing, walking, going downstairs, going upstairs, and biking) using six different types of mobile phones. (2) UCI (Reyes-Ortiz et al., 2016) captures six activities (i.e., standing, sitting, lying, walking, going downstairs, and going upstairs). (3) Motion (Malekzadeh et al., 2019) involves 24 participants performing six activities (sitting, standing, walking, going upstairs, going downstairs, and jogging). (4) Shoaib (Shoaib et al., 2014) includes 10 participants engaged in seven activities (sitting, standing, walking, going upstairs, going downstairs, jogging, and biking). In addition, we also select datasets (i.e., PAMAP2 (Reiss, 2012)) with more complex activities to show the performance of LanHAR. PAMAP2 consists of data collected from 9 participants performing 12 different activities, including both simple activities such as sitting, standing, and walking, and more complex activities such as vacuum cleaning and ironing.
@@ 4.1.2. Data preprocessing
We preprocess all datasets to ensure consistent sampling rates and window sizes. The sampling rate is set to 20 Hz to simulate a resource-constrained scenario, following prior studies (Xu et al., 2023). For cross-dataset activity recognition, we focus on four common activities (walking, going upstairs, going downstairs, and sitting) across the four datasets. We select one dataset as the source and one of the remaining three as the target, resulting in 12 combinations in total. Datasets are categorized based on their roles as either source or target. For source datasets, each is split into a training set (80%) and a validation set (20%). For target datasets, the entire dataset is used as the test set to directly assess cross-dataset performance. For new activity recognition, we train the model on the four common activities from the source dataset. We then select activities from the target dataset that differ from the four common activities as new activities. Finally, we evaluate the model’s ability to recognize these new activities by determining which activity labels the new data correspond to.
4.1.3. Baselines
We compare our method LanHAR with three categories of baselines: conventional HAR, cross-dataset HAR, and new activity baselines.
(i) Conventional HAR baselines: DCNN (Yang et al., 2015) is widely used in the field of human activity recognition for their ability to automatically extract local features from time-series sensor data. Transformers (Vaswani, 2017) utilize self-attention mechanisms to capture long-range dependencies in sensor data. LIMU-BERT (Xu et al., 2021) leverages masked language modeling and self-supervised learning to effectively utilize unlabeled IMU sensor data.
(ii) Cross-dataset HAR baselines: SDMix (Lu et al., 2022) It is a method for cross-domain activity recognition that uses the mixup method that addresses semantic inconsistencies caused by domain gaps. UDAHAR (Chang et al., 2020) is a method for cross-dataset activity recognition using domain adaptation. UniHAR (Xu et al., 2023) is a method for cross-dataset activity recognition. It designs physical-informed augmentation methods to improve the cross-dataset HAR performance. UniMTS (Zhang et al., 2024) is a unified pre-training framework designed for motion time-series data.
(iii) New activity baselines: DCNN (Yang et al., 2015) To adapt it for recognizing new activities, we fine-tune it using a small portion of data from the target dataset. Nuactiv (Cheng et al., 2013) is a method designed for new activities. It predefined activity label-related attributes: actions of a specific part of the body. UniMTS (Zhang et al., 2024) is a unified pre-training framework designed for motion time-series data, with zero-shot and few-shot generalization across diverse human activity recognition tasks. HARGPT (Ji et al., 2024) leverages LLMs as zero-shot human activity recognizers by framing sensor-based activity recognition as a natural language processing task.
(iv) Variants of our model: (1) LanHAR without Subtasks of Text Encoder (w/o STE). We remove the subtasks from the text encoder and focus its training on alignment. (2) LanHAR without Training of Text Encoder (w/o TTE). We remove the training part of the text encoder and directly use a pre-trained BERT model to replace it. (3) LanHAR with LLAMA2 (w LLAMA2). We replace GPT-4 with LLAMA2 to obtain semantic interpretations. (4) LanHAR with GPT-3 (w GPT-3). We replace GPT-4 with GPT-3 to obtain semantic interpretations. (5) LanHAR with Different Language Model (w DLM). We use the transformer architecture to replace BERT in LanHAR. (6) LanHAR without Iterative re-Generation (w/o IG). We remove the iterative re-generation method. (7) LanHAR without Iterative re-Generation and designed Prompt (w/o IGP). We remove the iterative re-generation method and the designed prompt. We directly use a simple prompt only containing data and task requirements.
4.1.4. Metric
We evaluate our model using two metrics: accuracy and F1 score. For cross-dataset HAR performance, we use accuracy to measure the correctness of the classification results for the four selected activities. For new activities, we use accuracy to evaluate the correctness of the classification results for the new activity only.
4.1.5. Implementation
We implement our method and baselines using PyTorch 1.9.0 in a Python 3.7 environment and train them on two NVIDIA RTX A5000 GPUs, each with 24 GB of memory. For LLMs, we choose GPT-4 as our LLMs to obtain semantic interpretations of sensor readings and activity labels. For the text encoder, we use the BERT-base architecture as its structure and initialize it with the pre-trained BERT-base parameters (Devlin et al., 2018). For the sensor encoder, we set the embedding dimension to $768$ to maintain consistency with pre-trained BERT. The multi-head attention parameter is set to $2$ and the encoder layer consists of $3$ layers. We utilize the AdamW optimizer with a learning rate of 1e-5 and set the batch size to 32. For iterative re-generation, we set the threshold for intra-cluster distance change to 0.0005.
4.2. Overall performance
We compare our method on various baselines, and the results are presented in Table 2. Overall, our method demonstrates superior average performance across 12 different source and target dataset combinations, leading in the majority of settings. Compared to conventional HAR baselines, our method consistently outperforms them. These traditional baselines are not specifically designed for cross-dataset HAR, and their strong performance typically depends on large, carefully curated datasets. Compared to cross-dataset HAR baselines, our method outperforms them (except in the setting from Shoaib to Motion). Although SDMix and UDAHAR are also designed for cross-dataset HAR, they have specific requirements. For instance, UDAHAR may be more suitable for datasets with smaller distributional gaps. This consistent performance demonstrates the robustness and reliability of our method in diverse conditions and datasets.
@@ 4.2. Overall performance
LanHAR Accuracy 0.830 0.812 0.730 0.855 0.811 0.782 0.697 0.764 0.730 0.717 0.789 0.768 0.774
F1 scores 0.804 0.795 0.693 0.808 0.765 0.718 0.617 0.713 0.724 0.687 0.730 0.721 0.731
@@ 4.3. New activity performance
We compare our method with four new activity recognition baselines. Specifically, we train the model on four common activities from the source dataset and evaluate it on different activities from the target dataset that are not among the four common activities, treating them as new activities. The trained model is then used to recognize these new activities, i.e., determining which activity new data belong to. Fig. 10 presents the performance of all methods in recognizing new activities across multiple target datasets.
For new activity recognition, our model outperforms all the baselines: Our model performance is superior to that of a fine-tuned DCNN model, which indicates its advantage over simple fine-tuning on small datasets. Additionally, our model performs better than Nuactiv, which relies on predefined attributes for activities. Defining these attributes is challenging and varies significantly across different activity contexts, directly impacting Nuactiv’s performance. This limitation hinders Nuactiv’s performance in cross-dataset settings. Similarly, HARGPT, as a prompt-based HAR method, has limited ability to recognize new activities. UniMTS, although trained on large-scale simulated data, still does not perform well in zero-shot scenarios on real-world datasets.
Figure 9. New activity performance
Figure 10. Complex activity performance
4.4. Complex activity performance
In addition to simple activities such as sitting and walking, we also select datasets containing more complex activities such as vacuum cleaning, ironing, and rope jumping to validate the effectiveness of our model. However, since different datasets contain different complex activities, to our knowledge, there are no two truly comparable datasets for cross-dataset HAR evaluation. Therefore, instead of using two different datasets, we select a single dataset and split it based on individuals to simulate cross-dataset performance. Specifically, we randomly partition the PAMAP2 dataset into two subsets based on individuals’ IDs and select the partition with the highest KL divergence, indicating the greatest data variance across two subsets, as the final partitioning result. The evaluation result is presented in Fig. 10. As illustrated, our model demonstrates superior performance compared to other models on datasets that include complex activities, achieving an improvement in accuracy of at least 4.58% and at most 13.78%, and an increase in F1 scores of at least 5.43% and at most 11.32%.
@@ 4.5. Category-level HAR performance
We further evaluate the category-level accuracy of HAR performance. An activity is considered correctly classified if it is assigned to a similar activity category and incorrectly classified if it falls into a dissimilar category. The motivation for this evaluation is that even if a model does not provide the correct prediction, we expect its misclassification to be a similar activity rather than a completely unrelated one. Specifically, we group seven activities (walking, jogging, biking, sitting, standing, going upstairs, going downstairs, and lying) into three categories based on their similarity: Category 1 (walking, jogging, biking), Category 2 (sitting, standing, lying), and Category 3 (going upstairs, going downstairs). We compare our method against all the baselines. The average results are shown in Fig. 12. As shown, our method improves category-level HAR accuracy compared to the baselines. This improvement has the potential to reduce the impact of classification errors on downstream applications that require a broader activity categorization rather than fine-grained recognition.
@@ 4.6.1. The effect text encoder
To assess the impact of the subtasks in the text encoder, we compare our method with a variant (i.e., w/o STE), as shown in Fig. 12. In the w/o STE variant, we remove the subtasks (i.e., the two contrastive learning and reconstruction tasks) and focus only on aligning the semantic interpretations of sensor readings and activity labels. As shown in the figure, removing the subtasks leads to a decline in performance compared to our full model. This demonstrates that the subtasks we designed help the model better understand IMU sensor readings and HAR-related semantics.
To further explore the effect of the text encoder on semantic interpretation alignment, we replace the carefully designed text encoder with a pre-trained BERT model and directly use it for training in Stage 2. The results are shown in Fig. 12. From the figure, it is clear that the performance of w/o TTE is lower than both w/o STE and our full model. This suggests that pre-trained language models without fine-tuning cannot fully capture HAR-related semantics or effectively align different HAR-related semantic interpretations.
@@
4.6.2. The impact of different pre-trained LMs
Our text encoder is flexible and allows the integration of different language models. In our work, we employ a BERT-based architecture, fine-tuned on our designed tasks. To investigate the impact of different language model architectures on HAR performance, we replace the BERT-based encoder with GPT-2. As shown in Fig. 14, this replacement results in decreased HAR performance. BERT is generally more suitable for tasks that require understanding sentence meaning and comparing text similarities, as its architecture and pre-training are optimized for capturing semantics and contextual relationships in a bidirectional manner. In contrast, GPT-2, with its unidirectional processing, is less effective at understanding sentence-level meaning and comparing text similarity, though it excels in generating coherent and contextually relevant text. This explains the performance drop when GPT-2 is used for HAR tasks.
Figure 13. The impact of different pre-trained Language Models as text encoders
Figure 14. The impact of different Large Language Models
@@ 4.6.3. The impact of different LLMs
Given the varying capabilities of different LLMs, we explore how three LLMs (GPT-4, GPT-3, and LLAMA2) impact the interpretation of sensor readings and activity labels in HAR performance. As shown in Fig. 14, the performance of GPT-3 and LLAMA2 is noticeably lower than that of GPT-4, highlighting the importance of the LLM’s capability in language-centered human activity recognition. These results suggest that stronger LLMs with advanced reasoning abilities are better suited for enhancing HAR performance.
4.6.4. The impact of different response qualities of LLMs
To ensure the quality of the LLM’s interpretation of sensor readings, we (i) design specific prompts to guide the LLM’s responses and (ii) develop an iterative re-generation method to filter out low-quality semantic interpretations and generate improved ones. To demonstrate the effectiveness of the automatic iterative re-generation method, we compare the relative KL divergence of the same activities across different datasets before and after applying the method. Specifically, for each activity, we measure the relative KL divergence between any two datasets both with and without applying the iterative re-generation method (i.e., “w iterative re-generation” and “w/o iterative re-generation”), as shown in Fig. 16. After applying the method, the average relative KL divergence for the same activity across different datasets decreases by 16.92%, indicating that the automatic iterative re-generation method significantly improves the response qualities of LLMs and reduces the distribution gap in heterogeneous data.
Figure 15. Relative KL divergence with and without the iterative re-generation method
@@ 4.6.4. The impact of different response qualities of LLMs
To verify the impact of these methods on HAR performance, we compare our method with two variants: one using only simple prompts (w/o IGP) and another using only specifically designed prompts (w/o IG). As shown in Fig. 16, the performance when using only simple prompts is the lowest, indicating that LLMs require guidance to generate high-quality responses. The combination of specifically designed prompts and the automatic iterative re-generation method performs the best, demonstrating that our method effectively enhances the quality of LLM responses, leading to improved performance in human activity recognition.
@@ 4.7. Parameter sensitivity
In our model, selecting the appropriate parameters for the sensor encoder is crucial, as it directly affects its suitability for deployment on mobile devices (e.g., ensuring the parameter size is sufficiently small). In this section, we analyze the impact of the embedding dimensions and the number of encoding layers on performance.
@@ 4.7.1. Effect of different number of attention heads
Fig. 18 shows the performance comparison $w.r.t.$ the number of attention heads in the sensor encoder. Overall, the performance is relatively stable under different numbers of attention heads. We choose the number of attention heads of 2 because it provides good performance while reducing the model’s parameter size.
@@ 4.7.2. Effect of different number of encoding layers
Fig. 18 shows the performance comparison with respect to different numbers of encoding layers in the sensor encoder. As observed, a smaller number of encoding layers (e.g., 2 layers) results in insufficient representation, whereas a larger number (e.g., 6 layers) increases model complexity. We select 3 encoding layers as they strike a balance, delivering good performance without significantly increasing model complexity.
@@
4.8. Deployment analysis
To evaluate the suitability of the sensor encoder for mobile deployment, we analyze its parameter size and inference time. Deploying LanHAR on mobile devices involves three key steps: (i) developing an Android application, (ii) converting the model into a .pt file, and (iii) integrating it for processing accelerometer and gyroscope data for activity recognition. For our case study, we select the Pixel 7 smartphone as the deployment platform. The sensor encoder, combined with the activity label embeddings, occupies approximately 40 MB of memory, of which 0.01 MB is allocated for the embeddings. To assess its real-time performance, we conduct 500 inference trials, measuring the time required to recognize an activity. The results indicate an average processing time of 0.027 seconds, with the minimum and maximum times recorded at 0.037 seconds and 0.023 seconds, respectively. For comparison, we evaluate the model’s inference efficiency on a single A5000 GPU with a batch size of 1. The average inference time for a single activity is 0.0007 seconds, with the minimum and maximum times recorded at 0.0006 seconds and 0.001 seconds, respectively. Although inference on the smartphone is significantly slower than on the GPU, it remains within a practical range for real-world mobile applications, supporting real-time inference with an IMU sampling frequency of around 40 Hz (i.e., 0.025 seconds per reading). This demonstrates that LanHAR can operate effectively on mobile devices. To further enhance efficiency, future optimizations, such as quantization and hardware acceleration, could be explored.